# Making Differential Privacy Easier to Use for Data Controllers and Data Analysts using a Privacy Risk Indicator and an Escrow-Based Platform

### v1

@@ 3.2 The Find-$\epsilon$-from-PRI Algorithm
1 /* Initialize set of $\epsilon$ values to test; minimum value is $\epsilon_{c}$/
2 $\mathcal{E}\leftarrow$ initialize-eps-to-test($\epsilon_{c}$)
3 $x^{p}\leftarrow\textsc{project-query-attrs}(x,q)$
4 /* Compute $q(x_{-i})$ for all $i$ using memoization */
5 $C\leftarrow map()$, $L\leftarrow list()$
6 for $x^{p}_{i}\in x^{p}$ do
7       if $x^{p}_{i}\notin C$ then
8             $C[x^{p}_{i}]\leftarrow q(x^{p}_{-i})$
9      $L.add(C[x^{p}_{i}])$
10for $\epsilon\in\mathcal{E}$ do
@@ 8             $C[x^{p}_{i}]\leftarrow q(x^{p}_{-i})$
12       $r\leftarrow\mathcal{M}(x,q,\epsilon)$
13       /* Compute PRI for each individual */
14       $PRI\leftarrow list()$
15       for $l\in L$ do
16             $PRI_{i}\leftarrow\lVert r-l)\rVert_{1}$
17             $PRI.add(PRI_{i})$
18      $PRI_{max},PRI_{p}\leftarrow\max(PRI)$, get-percentile$(PRI,p)$
19       /* A suitable $\epsilon$ is found when the ratio is close to 1 */
20       if $PRI_{max}/PRI_{p}<1+\tau$ then
21             $\epsilon_{c}\leftarrow\epsilon_{c}+\epsilon$
@@
Algorithm 1 shows the pseudo-code for finding $\epsilon$ from PRI and releasing DP output. First, the algorithm initializes the set of $\epsilon$ to choose from (line 1). Although the number of possible $\epsilon$ can be infinite, to ensure the algorithm runs in bounded time, we set a range (e.g., from $\epsilon_{c}$ to 10) and include representative $\epsilon$ value in fixed intervals (e.g., 0.1). The lower bound $\epsilon_{c}$ starts at 0 and updates for each answered query (we explain how to handle multiple queries in Section 3.3). The algorithm then projects relevant query attributes from the dataset (line 1) and computes the query output for each neighboring dataset (line 1 - 1). Next, the algorithm iterates over the list of $\epsilon$ from largest to smallest (line 1) and stops when it finds the first suitable value. For each $\epsilon$ it tests, the algorithm computes the current query output $r$ by applying the given DP mechanism $\mathcal{M}$ with parameter $\epsilon$ (line 1), and use it to compute the corresponding PRI for each individual (line 1 - 1). Finally, the algorithm computes the ratio of the maximum PRI over the PRI at the given $p$th percentile (line 1). If the ratio is close to 1 within the threshold $\tau$ (we use 0.01 in our implementation), then a suitable $\epsilon$ has been found. The algorithm updates the lower bound $\epsilon_{c}$ and returns the DP output $r$ (line 1 - 1).
Algorithm data flow. The algorithm is used by controllers to find $\epsilon$ from the PRIs and release DP output. The controller runs the algorithm on their input dataset and query. The algorithm first computes the PRIs, finds a suitable $\epsilon$ based on the controller’s privacy requirement, and returns the DP output computed using this $\epsilon$. Nothing is released by default to analysts. The PRIs are a by-product of which the sole purpose is to help the controller find $\epsilon$, and they are never released to the data analysts (we will discuss how such promise is enforced in Section 4.3. The $\epsilon$ value may need to be released on certain occasions; since finding $\epsilon$ is a post-processing step to computing PRI, which is DP, its release is also DP.
DP guarantee. The algorithm satisfies $\epsilon$-differential privacy. The algorithm only returns the query output $r$ when a suitable $\epsilon$ can be found. The output $r$ is computed using a standard $\epsilon$-differentially private mechanism (i.e., Laplace Mechanism); thus, the algorithm shares the same DP guarantee. This is no different than a standard DP deployment, only that $\epsilon$ is chosen automatically instead of manually by the controller.
3.3 Bounding Privacy Loss Across Queries
@@
The challenge of running multiple queries. In practice, analysts often want to run multiple queries. Combining these queries’ output will reveal new information, eventually breaking privacy as per the fundamental Law of Information Recovery [22]. In today’s deployment of DP, the data controller needs to set a total privacy loss budget a priori on the dataset and stop accepting queries when the cumulative $\epsilon$ exceeds the budget. While the total budget provides a rigorous bound of privacy loss in the context of DP, it abstracts away from the actual privacy outcomes, and setting it without context is extremely challenging [23]. To further exacerbate the problem, the data analyst needs to decide the queries they want to run a priori since they need to decide how to utilize the total budget for their queries. This poses a major obstacle for analysts who do not have a pre-determined set of queries.
@@
The solution. Instead of quantifying the maximum acceptable privacy loss, which is hard, we keep track of the already consumed budget on a dataset, i.e., the sum of $\epsilon$ of all previously answered queries; we call it $\epsilon_{c}$. For an incoming query, when using Algorithm 1 to find $\epsilon$, we truncate the range of $\epsilon$ values to consider only those larger than $\epsilon_{c}$ (line 1). We reject answering this query if we fail to find a suitable $\epsilon$ that is larger than $\epsilon_{c}$, meaning the PRIs cannot offer adequate privacy protection for any possible $\epsilon$.
Our solution creates a dynamic bound of total privacy loss. As the cumulative privacy loss, $\epsilon_{c}$, keeps growing for each answered query, it becomes more difficult to find a suitable $\epsilon$ that is larger than $\epsilon_{c}$ since less noise can be added to make the PRIs meet the controller’s requirement. At each point that we fail to find a suitable $\epsilon$ for a query, the total privacy loss is bounded. Using DP’s sequential composition property, releasing the output of those answered queries satisfies $\epsilon_{c}$-differential privacy. The bound is refreshed if a suitable $\epsilon$ is found for a different query. This resulting dynamic bound helps controllers bound the total privacy loss without the need to set the total budget a priori. Since the algorithm automatically finds $\epsilon$ (or rejects queries) based on this dynamic bound, it also removes the need for the analyst to decide how to distribute their privacy budget across different queries a priori.
3.4 Discussion and Extensions
Algorithm complexity and optimizations. The algorithm has a worst-case time complexity of $O(nm)$, where $n$ is the number of records in the dataset, and $m$ is the number of $\epsilon$ values we test. Computing the PRI of each individual requires computing the query output of each neighboring dataset (line 1 - 1). When the number of individual records is high, this becomes a performance bottleneck: using memoization saves computation if the number of unique projected records is smaller than $n$. In addition, the query computation can be parallelized since each computation is independent of the other. We implement multi-processing and compute the query output for a chunk of records in each process, and we show in Section 5.3.1 that the algorithm is both efficient and scalable.
Expressing privacy preferences. A modification of our algorithm permits the controller to specify protection preferences over certain groups of individuals whose privacy should be protected similarly. For example, they can specify individuals whose ages are above 60 and those younger need to receive the same privacy protection. This feature can easily be incorporated into the algorithm by letting the controller change the ratio to compute. In the example, the controller can choose to compare the median PRI of the two age groups and choose the $\epsilon$ that makes the ratio of the two medians close to 1.
Communicating Privacy-Accuracy Trade-off. In some circumstances, the data controller may want to consider the accuracy of the released output when deciding on the level of privacy protection they want to offer each individual. Prior work such as DPComp[28] and ViP[40] have explored different ways to visualize privacy-accuracy trade-off in differential privacy. The PRIs can complement the existing interfaces by providing more information on the relationship between output accuracy and individual privacy implications.
4 Helping Analysts Interpret Utility
In this section, we first propose a simple utility signaling protocol to help analysts understand the impact of DP on their downstream applications to address Challenge 2 (Section 4.1). Then, we assemble a set of requirements for a DP platform, which we derive from the solutions proposed (Section 4.2). In Section 4.3, we introduce a new implementation of our new algorithm and protocol on top of a data escrow that meets those requirements.
4.1 The utility signaling protocol
Today’s DP deployments. In the central DP model, we differentiate among individuals who own the original data, controller that manages individuals’ data on their behalf, platform the infrastructure where computation takes place, and analyst the actor demanding computation over individuals’ data. The analyst submit queries to a platform operated by a controller. The platform computes query output using DP and releases them to the analyst with the controller’s approval.
We consider the case where the analyst has a downstream task with an associated evaluation function, g, that returns the utility of the downstream task. The function takes the set of query output, S, that the analyst received as input. Following the notation used in Section 2.2), $S=\{\mathcal{M}(x,q,\epsilon)\mid x\in\mathcal{X}^{n}\}$, $q:\mathcal{X}^{n}\rightarrow\mathbb{R}^{k}$. For example, an analyst queries census data at the county level (i.e., the US Census Bureau is the controller and citizens are the individuals). The set of output $S$ are aggregate statistics on population count, education level, housing prices, etc. The analyst’s task is to build a machine learning model to predict the median income of each county based on those statistics, and the evaluation function returns the prediction accuracy, which is g(S). Since the queries are computed using DP, releasing $g(S)$ is also DP by post-processing.
The problem. The controller uses DP to release noisy output to protect the privacy of individuals, but the analyst does not know how much such noise affects their downstream task and whether the output is useful. In the example, an analyst who obtains a prediction accuracy of $88\%$ does not know how much the accuracy would have been without noise; thus, they cannot decide whether it is worth investing more effort in improving the accuracy of a model. Clearly, the platform cannot tell the analyst how much noise was added because that would jeopardize the privacy guarantee.
A proposed solution: a utility signaling protocol. We propose a protocol to help analysts interpret the implications of the added noise on the quality of their downstream task by letting them compare to the quality they would have obtained over the ground truth output. Let $S^{\prime}=\{q(x)\mid x\in\mathcal{X}^{n}\}$ be the set of ground truth output (i.e., without applying DP). The protocol allows the analyst to obtain both g(S) and g(S’), thus signaling the utility loss due to DP.
Sending the ground truth output to the analyst so they run their downstream task is not feasible as it breaks the premise of protecting the data in the first place. Instead, the analyst sends their downstream task to the platform. The platform invokes the evaluation function $g$ over both the set of noisy output $S$ and the set of ground truth output $S^{\prime}$. Then, if the controller approves, both $g(S)$ and $g(S^{\prime})$ are released to the analyst to signal the loss in utility due to DP. In the example, the platform trains a model over both the DP census statistics and the ground truth statistics, then applies the evaluation function the analyst shared, obtaining $88\%$ and $90\%$ accuracy, respectively. Upon receipt of both numbers, the analyst understands the utility effect of DP and can decide whether it is worth investing more effort in improving the model.
Assessing Privacy Implications of Releasing $g(S^{\prime})$: Any information release may leak information about the individuals’ data, e.g., even the accuracy of a model (consider the trivial case where $S^{\prime}$ contains only one value, the analyst can simply use the evaluation function to test different values until they obtain an accuracy of 100%). Since DP is not being applied when releasing $g(S^{\prime})$, it is ultimately at the discretion of the controller to decide whether or not to release $g(S^{\prime})$. In the example, the controller may be comfortable releasing the accuracy of the machine learning model trained on the ground truth census statistics (in addition to the model accuracy on the DP statistics), but not releasing the model itself. Our utility signaling protocol does not attempt to address the possibility of various attacks the analyst may perform to infer individual data from $g(S^{\prime})$. Instead, the escrow-based platform we describe next allows the controller to examine the analyst’s downstream task and associated evaluation function and ensures $g(S^{\prime})$ will not be released without the controller’s permission. For many applications, the benefit of allowing the analyst to observe $g(S^{\prime})$ can be tremendous.
4.2 Platform Requirements
We propose a new platform to support the Find-$\epsilon$-from-PRI algorithm that helps controllers (addressing Challenge 1) and the utility signaling protocol that helps analysts (addressing Challenge 2). Such a platform must meet the following requirements:
  • R1. Control Data Flow: Analysts should never see the raw data or PRIs. And they should not submit queries and downstream tasks over the raw data directly without the controller’s approval (e.g., the controller may approve releasing the model accuracy to signal utility loss).
  • R2. Trust: The new protocol requires more trust on the platform. Trust is a broad concept, and its meaning varies depending on context[7]. In the context of DP, trust means the platform adheres to prespecified behavior and the DP mechanism upholds its privacy guarantee[7][43]. Individuals must trust the controller will protect their data in a responsible manner (i.e., using DP). Controllers must trust the platform keeps the data secure and runs computations on the data correctly. With the new utility signaling protocol, analysts must also trust the platform runs their downstream task correctly. We want to reduce the trust those entities have on the platform.
  • R3. High Performance: The platform must limit the overhead, which is at odds with the pressures of the new Find-$\epsilon$-from-PRI algorithm that helps controllers at the cost of higher computational overhead.
We highlight a tension between R2 and R3. Improving trust often requires employing techniques (e.g., cryptographic primitives) that increase the computational overhead, potentially leading to impractical solutions. We now explain the design of a platform that addresses these three requirements.
4.3 Escrow-Based DP Platform
We propose a new platform, based on an escrow architecture similar to the proposed in  [50], that lets controllers maintain fine-grained control of data flow (meeting R1), reduces the trust on the controller and platform by both individuals and analysts (meeting R2) all while maintaining high performance to meet R3.
4.3.1 Controlling Data Flow to Implement the utility signaling protocol
A data escrow supports delegated computation, which means the escrow runs computation on behalf of the computation owner. Implementing a platform with an escrow lets analysts delegate their downstream tasks to the platform, thus obtaining the two utilities $g(S)$ and $g(S^{\prime})$. We present the end-to-end protocol in Figure 3, which consists of seven steps.
Figure 3: End-to-end utility signaling protocol. We illustrate the protocol in the central DP deployment scenario where the controller has access to the raw data.
1. Registering data. The controller registers the dataset containing individual records with the escrow.
2. Registering application. The controller registers an application that uses DP to run queries on the registered dataset, which includes the Find-$\epsilon$-from-PRI algorithm. Any derived data product, such as the PRI, will not be released without the controller’s permission. The PRI can also be computed separately as an independent function inside the application if the controller wants to use it for other purposes.
3. Creating policy. The data controller creates a policy to let the analyst run queries on the registered dataset using the registered application; a policy in the escrow indicates who can run what computation on what data.
4. Issuing Queries. The analyst issues queries to the application. The application answers each query and stops when a suitable $\epsilon$ can no longer be found. After checking that the query computation only ran on data accessible to the analyst, the escrow releases the set of DP output $S$ to the analyst because the controller wrote the corresponding policy. The set of ground truth output $S^{\prime}$ is registered inside the escrow since it is a derived data product of the computation.
Up to this point, the protocol provides analysts with the same information they can access in a typical central DP deployment. Now, we extend the protocol to implement the utility signaling component to help analysts understand the impact of DP on their downstream tasks.
5. Registering downstream task. The analyst registers their downstream task and an associated evaluation function $g$ as an application in the escrow. The evaluation function is the only function within the application that is available for the escrow to invoke and anyone to create policy on; if exists, the rest of the functions should be invoked from the evaluation function to perform other necessary computations. The escrow ensures that only the evaluation function’s output (and nothing else) will be released to the analyst.
6. Creating policy for downstream task. After inspecting the application submitted by the analyst and determining it is acceptable to release $g(S^{\prime})$, the controller creates a policy to run the evaluation function $g$ on the ground truth output $S^{\prime}$.
7. Evaluating downstream task. With the controller’s permission, the analyst obtains both $g(S)$ and $g(S^{\prime})$. By inspecting the difference between the two utilities, the analyst learns the impact of DP on their task, thus helping them decide whether the DP output is sufficiently good for their task.
By exploiting the delegated computation features of the data escrow, the platform controls data flow, which is necessary to implement the utility signaling protocol.
4.3.2 From Trusted to Trustworthy, while Fast
A trusted platform adheres to prespecified behaviors[7], in the context of DP this means an adversary only has access to the output of a DP mechanism, but not any of its internal state and any raw or derived data[22]. This is infeasible to achieve in practice. A central platform, as in the central DP model, can claim to be trusted, but such claim does not preclude it from being compromised and intentionally leaking data in the clear, thus breaking the DP guarantee[43]. A trustworthy platform, on the other hand, implements different techniques to enforce the promise that not only will it never behave unexpectedly, it will not pose harm (to the users) in any way[27]. An escrow-based platform reduces the impractical assumptions of a trusted platform and supports trustworthy computation by ensuring:
  • No data or derived data product (e.g., query output) ever leaves the platform without explicit permission from the controller, and only explicitly allowed computation runs on existing data; this is the basis of the control of data flow.
  • Data is encrypted end-to-end, even during computation. End-to-end encryption can be achieved in many ways; for instance, DP platforms leveraging secure multiparty computation[42][9][10][49][45], homomorphic encryption[4][46], Oblivious RAM[12] and trusted execution environment[44]. The prototype we use for our evaluation uses AMD Secure Encrypted Virtualization(SEV)[33].
  • Tasks that are sent to the escrow are the tasks that run. The escrow is registered with an escrow application; the platform ensures that the correct version of the application is running using techniques such as remote attestation[15] or formal verification[38][35]. Our prototype uses hardware-rooted remote attestation enabled by AMD SEV.
Figure 4: Left: central platform. Right: escrow-based platform
By computationally enforcing the promises above, the escrow reduces the trust that both individuals and analysts must place in the platform while maintaining the same accuracy guarantee as the central DP model. Furthermore, the escrow application can be written so that controllers do not gain access to the raw data but only the PRIs (i.e., by writing a policy that only reveals the PRIs to the controller), thus illustrating multiple escrow-based deployment scenarios. Consider Figure 4, which shows how individuals, controller, and analyst interact with the central platform in CDP and the escrow-based platform. In CDP, the controller manages individuals’ data using the central platform and sets $\epsilon$ manually; the individuals need to trust both the controller to protect their privacy (i.e., by choosing a good $\epsilon$) and the platform to operate correctly without leaking their data. Whereas using the escrow-based platform, individuals send data directly to the platform, and the controller only has access to PRIs computed by Algorithm 1. The escrow invokes Algorithm 1 to find $\epsilon$ on the controller’s behalf and releases the output to the analyst, who delegates their downstream task to the escrow to obtain utility.
5 Evaluation
Our evaluation shows our solutions address both Challenge 1 to help controllers, and Challenge 2 to help analysts. We answer the following research questions for our evaluation:
  • RQ1: Does presenting the privacy risk indicator help data controllers choose $\epsilon$?

### v2

@@ 3.2. The Find-$\epsilon$-from-PRI Algorithm
1 /* Initialize set of $\epsilon$ values to test */
2 $\mathcal{E}\leftarrow$ initialize-eps-to-test()
3 /* Compute $q(x_{-i})$ for all individual $i$ using memoization */
4 $C\leftarrow map()$
5 for $x_{i}\in x$ do
6       if $x_{i}\notin C$ then
7             $C[x_{i}]\leftarrow q(x_{-i})$
8
9 for $\epsilon\in\mathcal{E}$ do
@@ 7             $C[x_{i}]\leftarrow q(x_{-i})$
11       $r\leftarrow\mathcal{M}_{L}(x,q,\epsilon)$
12       /* Compute corresponding PRI of each individual */
13       $PRIs\leftarrow list()$
14       for $x_{i}\in C$ do
15             $PRI_{i}\leftarrow\lVert r-C[x_{i}]\rVert_{1}$
16             $PRIs.add(PRI_{i})$
17      $PRI_{min},PRI_{max}\leftarrow\min(PRIs),\max(PRIs)$
18       /* Check whether $\epsilon$ satisfies controller’s privacy preference */
19       if $PRI_{min}/PRI_{max}\geq 1-(p/100)$ then
@@ 17      $PRI_{min},PRI_{max}\leftarrow\min(PRIs),\max(PRIs)$
Detailed Algorithm. Algorithm 1 shows the pseudo-code for finding $\epsilon$ from PRI and releasing DP output. First, the algorithm initializes the set of $\epsilon$ to choose from (line 1). Although the number of possible $\epsilon$ values is infinite, we set an upper bound and a range, e.g., 10 and 0.1, to ensure a finite set of options, e.g., 100 in this example. Then, for each $x$’s neighboring dataset $x_{-i}$, the algorithm computes $q(x_{-i})$ (line 1 - 1), which will be used to compute the PRI for each individual $i$. Next, the algorithm iterates over the set of $\epsilon$ in descending order (line 1). For each $\epsilon$, the algorithm computes the query output $r$ by applying the Laplace Mechanism $\mathcal{M}_{L}$ with parameter $\epsilon$ (line 1) and computes the corresponding PRI of each individual (line 1 - 1). Then, given the list of PRIs, the algorithm computes $PRI_{min}$ and $PRI_{max}$ (line 1). If $PRI_{min}/PRI_{max}\geq 1-(p/100)$, then the algorithm has found a suitable $\epsilon$ and a query output, so it stops and returns the output.
Figure 2. Example of $\epsilon$ chosen by the Find-$\epsilon$-from-PRI algorithm with different percentage $p$ over 50 runs
Running the algorithm with the same dataset $x$, query $q$ and parameter $p$ may identify different $\epsilon$ because the same $\epsilon$ may produce different PRI profiles across executions of the mechanism; simultaneously, similar PRI profiles may be produced by different $\epsilon$ values. What is important is that the final query output the analyst releases has been subject to DP and the privacy loss of the individuals is consistent with the controller’s privacy preferences as expressed by $p$. In addition, we show empirically in Section 6.3.2 that stricter privacy preferences (i.e., smaller $p$) generally lead to smaller $\epsilon$. Figure 2 shows a boxplot representing the distributions of $\epsilon$ chosen by the algorithm with different $p$ (5, 25, 50, 75, 95) over 50 runs; for demonstration, we choose one of the five queries (Q5) we have run. As shown in the figure, as $p$ increases, the algorithm tends to choose larger $\epsilon$. Setting $p=5$ means the minimum and maximum PRI needs to be very close (within 5%), thus the algorithm mainly finds $\epsilon$ that is less than 1; in contrast, setting $p=95$ means the maximum PRI can be as much as 95% larger than the minimum PRI, thus the $\epsilon$ found can be very large.
Figure 3. Dataflow of Find-$\epsilon$-from-PRI Algorithm
Dataflows and leakage. It is critically important to understand what information is released to whom. Figure 3 shows the involved dataflows in a scenario where a controller uses the Find-$\epsilon$-from-PRI algorithm. Individuals send their data to the controller; this dataflow remains unchanged from traditional central DP models (Near et al., 2021). The analyst submits a query to the controller. The controller expresses their privacy preferences in terms of $p$ (and not in terms of $\epsilon$) and runs the algorithm to find a suitable query output, which is an output produced by a standard DP mechanism (Laplace Mechanism in our work) with a value for $\epsilon$ selected by the algorithm. At this point, no information has been returned to the analyst yet. The PRI associated with each individual is never released to the analyst. The query output that satisfies the controller’s privacy preferences is only released with the controller’s approval. The value for $\epsilon$ that led to the query output is also not released by default.
Releasing $\epsilon$. Under certain scenarios (Dwork et al., 2019), $\epsilon$ may also need to be released. Since $\epsilon$ is chosen based on PRIs, which are data-dependent, the controller should release $\epsilon$ differential privately by incorporating techniques such as Exponential Mechanism (e.g., Report Noisy Max) (McSherry and Talwar, 2007) or Sparse Vector Technique (Dwork et al., 2009; Lyu et al., 2016) in the algorithm. Instead of choosing the largest $\epsilon$ that passes the PRI criterion, the algorithm will choose and return the first $\epsilon$ (which may or may not be the largest) that passes the PRI criterion while preserving DP.
3.3. Bounding Privacy Loss Across Queries
@@
In practice, analysts often want to run multiple queries. Combining these queries’ output will reveal new information, eventually breaking privacy as per the fundamental Law of Information Recovery (Dwork et al., 2014). In today’s deployment of DP, controllers need to set a total privacy budget on the dataset a priori and stop accepting queries when the cumulative $\epsilon$ exceeds the budget (Dwork et al., 2019). The need for setting a total budget poses a critical challenge for controllers:
Controllers do not know the privacy implications of total privacy budget. Similar to choosing $\epsilon$ for a single query, when deciding which total budget to set, controllers face the same challenge of not knowing the actual privacy implications on individuals after releasing multiple query outputs. Oftentimes, they resort to selecting the total budget arbitrarily or through trial-and-error (Dwork et al., 2019).
@@
The solution. Instead of quantifying the maximum acceptable privacy loss, which is hard, the algorithm keeps track of the already consumed budget on a dataset, i.e., the sum of $\epsilon$ of all previously answered queries; we call it $\epsilon_{c}$. For an incoming query, the algorithm truncates the range of $\epsilon$ values to consider only those larger than $\epsilon_{c}$ (line 1). The algorithm rejects the query if it fails to find a suitable $\epsilon$ larger than $\epsilon_{c}$, meaning the controller’s privacy preferences cannot be met for any possible $\epsilon$.
Our solution creates a dynamic bound of total privacy loss. As $\epsilon_{c}$ keeps growing for each answered query, it becomes more difficult to find a suitable $\epsilon$ that is larger than $\epsilon_{c}$ since larger $\epsilon$ may not produce PRIs that satisfy controllers’ privacy preferences. At each point the algorithm fails to find a suitable $\epsilon$ for a query, the total privacy loss is bounded. The bound is refreshed if a suitable $\epsilon$ is found for a different query. A potential concern is whether there is additional privacy leakage using this dynamic bound. In Section 3.4, we show the bound can be formalized as a privacy odometer in fully-adaptive composition (Rogers et al., 2016) and there is no leakage.
Considering analysts’ preferences over queries. A natural tension exists between analysts and controllers: analysts want more accurate output, whereas controllers want to protect individuals’ privacy by releasing noisy output with degrading accuracy. Because it gets increasingly harder to find a suitable $\epsilon$ after answering multiple queries as $\epsilon_{c}$ keeps increasing, analysts may need to prioritize queries of which they need more accurate output. One approach to incorporate analysts’ preferences when finding $\epsilon$ is to allow them to assign weights (i.e., priority scores) to their queries, then controllers can adjust the parameter $p$ in the algorithm based on analysts’ preferences, e.g., they can set a higher $p$ for high-priority queries so the algorithm can find larger $\epsilon$ with more accurate output. Nevertheless, solving related problems, such as how to communicate analysts’ preferences to controllers, and how to help controllers decide an appropriate $p$ that takes both analysts’ preferences and their own privacy preferences into account, are outside the scope of the paper but remain interesting directions to explore.
3.4. End-to-End DP Guarantee
Implications of choosing $\epsilon$ adaptively. First of all, if the algorithm has only answered one query, releasing the query output to the analyst satisfies DP and the total privacy loss is bounded by the $\epsilon$ that is chosen by the algorithm. However, if the algorithm answered multiple queries (for the same dataset), for each query $\epsilon$ is chosen adaptively because it is chosen based on previously answered queries — a suitable $\epsilon$ needs to be larger than $\epsilon_{c}$. Then the total privacy loss is not bounded by any fixed constant since the algorithm does not require setting a total privacy budget. This brings the question of whether releasing multiple query outputs computed using adaptively chosen $\epsilon$ causes additional privacy leakage.
Privacy odometer in fully-adaptive composition. To answer this question, we construct a privacy odometer (Rogers et al., 2016) for the algorithm, which provides a probabilistic running upper bound on the realized privacy loss after answering each query. The privacy odometer is proposed to provide privacy guarantees in fully-adaptive composition (Rogers et al., 2016), in which the privacy parameters are chosen adaptively as a function of previous interactions.
Theorem 3.4 in  (Rogers et al., 2016) has proved that there is no additional leakage to choose privacy parameters adaptively if the privacy bound is computed using sequential composition. Therefore, the privacy odometer for the Find-$\epsilon$-from-PRI algorithm under sequential composition (Dwork et al., 2014) can simply be defined as:
Definition 0 (Privacy Odometer).